Ncert Solutions Maths class 12th

Get insights from 2.5k questions on Ncert Solutions Maths class 12th, answered by students, alumni, and experts. You may also ask and answer any question you like about Ncert Solutions Maths class 12th

  • 2.5k

    Questions

  • 0

    Discussions

  • 0

    Active Users

  • 65

    Followers

Have a question about Ncert Solutions Maths class 12th? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Let  E1:the  event  that  the  item  is  manufactured  on  machine  A         E2:the  event  that  the  item  is  manufactured  on  machine  B         E3:the  event  that  the  item  is  manufactured  on  machine  CLet  H  be  the  event  that  the  selected  item  is  defective.  Using Bayes'  Theorm∴     P(E1)=50100,  P(E2)=30100  and  P(E3)=20100 P(H/E1)=2100,  P(H/E2)=2100  and  P(H/E3)=3100∴      P(E1/H)=P(E1).P(H/E1)P(E1).P(H/E1)+P(E2).P(H/E2)+P(E3).P(H/E3)                               =50100×210050100×2100+30100×2100+20100×3100=100100+60+60=100220=1022=511Hence,  the  required  probability  is  511.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Let  E1:the  event  that  a  person  has  TB         E2:the  event  that  a  person  does  not  have  TBLet  H  be  the  event  that  the  person  is  diagnosed  to  have  TB.  Bayes'  Theorm∴     P(E1)=11000=0.001,  P(E2)=1−11000=9991000=0.999 P(H/E1)=0.99,  P(H/E2)=0.001∴      P(E1/H)=P(E1).P(H/E1)P(E1).P(H/E1)+P(E2).P(H/E2)                               =0.001×0.990.001×0.99+0.999×0.001=0.990.99+0.999                               =0.9900.990+0.999=9901989=110221Hence,  the  required  probability  is  110221.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Let    E1  be  the  event  of  selecting  BagI           E2  be  the  event  of  selecting  BagIIand  E3  be  the  event  that  black  ball  is  selected∴       P(E1)=26=13  and  P(E2)=1−13=23  P(E3/E1)=37  and  P(E3/E2)=47∴         P(E3)=P(E1).P(E3/E1)+P(E2).P(E3/E2)                         =13.37+23.47=3+821=1121Hence,  the  required  probability  is  1121.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Let      E1:The  event  that  the  letter  comes  from  TATANAGARand     E2:The  event  that  the  letter  comes  from  CALCUTTAAlso,  E3:The  event  that  on  the  letter,  two    letters  TA  are  visible.∴P(E1)=12,  P(E2)=12  and  P(E3E1)=28  and  P(E3E2)=17[?  For  TATA  NAGAR,  the  two consecutive   letters  visible  are  TA,AT,TA,AN,NA,AG,GA,AR]∴  P(E3/E1)=28and  [For  CALCUTTA,  the  two consecutive   letters  visible  are  CA,AL,LC,CU,UT,TT  and  TA]So,  P(E3/E2)=17Now  using  Bayes'  Therorm,  we  have∴P(E1/E3)=P(E1).P(E3/E1)P(E1).P(E3/E1)+P(E2).P(E3/E2)                  =12.2812.28+12.17=1818+114=187+456=711Hence,  the  required  probability  is  711.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Given  that  A1:A2:A3=4:4:2∴  P(A1)=410,  P(A2)=410  and  P(A3)=210  whereA1,A2  and  A3  are  the  three  types  of  seeds.Let  E  be  the  events  that  a  seed    and  E¯  be  the  events  that  a  seed  does  not  ∴  P(EA1)=45100,  P(EA2)=60100  and  P(EA3)=35100and  P(E¯A1)=55100,  P(E¯A2)=40100  and  P(E¯A3)=65100(i)  P(E)=P(A1).P(EA1)+P(A2).P(EA2)+P(A3).P(EA3)                     =410.45100+410.60100+210.35100                     =1801000+2401000+701000=4901000=0.49(ii)  P(E¯/A3)=1−P(E/A3)=1−35100=65100=0.65(iii)  ,  Bayes'  Theorem,  we  get         P(A2/E¯)=P(A2).P(E¯/A2)P(A1).P(E¯/A1)+P(A2).P(E¯/A2)+P(A3).P(E¯/A3)                              =410.40100410.55100+410.40100+210.65100=16010002201000+1601000+1301000                              =160220+160+130=160510=1651=0.314Hence,  the  required  probability  is  1651  or  0.314

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

  Referring to Exercise  Q.41,  we  will  use  here,  Bayes'  Theorem(i)  P(E2/F)=P(E2).P(F/E2)P(E1).P(F/E1)+P(E2).P(F/E2)+P(E3).P(F/E3)                             =26.1316.0+26.13+36.1=218218+36=218×1811=211(ii)  P(E3/F)=P(E3).P(F/E3)P(E1).P(F/E1)+P(E2).P(F/E2)+P(E3).P(F/E3)                              =36.116.0+26.13+36.1=36218+36=36×1811=911Hence,  the  required  probabilities  are  211  and  911

Thumbs Up IconUpvote Thumbs Down Icon